### v3

# CAOTE: KV Cache Eviction for LLMs via Attention Output Error-Based Token Selection

Abstract
While long context support of large language models has extended their abilities, it also incurs challenges in memory and compute which becomes crucial bottlenecks in resource-restricted devices. Token eviction, a widely adopted post-training methodology designed to alleviate the bottlenecks by evicting less important tokens from the cache, typically uses attention scores as proxy metrics for token importance. However, one major limitation of attention score as a token-wise importance metrics is that it lacks the information about contribution of tokens to the attention output. In this paper, we propose a simple eviction criterion based on the contribution of cached tokens to attention outputs. Our method, CAOTE, optimizes for error due to token eviction, by seamlessly integrating attention scores and value vectors. This is the first method to use information from the value vector on top of attention-based eviction scores. Additionally, CAOTE can act as a meta-heuristic method with flexible usage with any token eviction method. We show that CAOTE, when combined with state-of-the-art attention score-based methods, always improves accuracies on the downstream task for LLAMA3 and QWEN2.5 model families, indicating the importance of leveraging information from values during token eviction process.
1 Introduction
Large Language Models (LLMs) represent a large step forward in natural language processing by demonstrating remarkable proficiency in tasks such as text generation (Coenen et al., 2021), machine translation (Xiao et al., 2023), and question-answering (Robinson et al., 2022). Many of these tasks require handling long prompt inputs efficiently, particularly in retrieval-augmented generation (RAG), long-form document understanding (Liao et al., 2024), summarization (Zhang et al., 2024a), and multi-turn dialogue systems (Thoppilan et al., 2022)—collectively referred to as long-context LLMs. A key challenge in long-context applications is the increased latency during inference and generation, primarily due to the quadratic computational complexity of self-attention and the growing memory demands of handling long sequences.
To alleviate computational overhead, Key-Value (KV) caching is a widely adopted technique that enables faster inference in LLMs. It does so by storing the key-value states of previously processed tokens and reusing them when generating new tokens, reducing redundant computations (Pope et al., 2023). However, while KV caching significantly enhances efficiency, it comes at the cost of substantial memory consumption as mentioned in (Zhang et al., 2024b) , especially in long-context setups. In such cases, the memory footprint of the KV cache often exceeds the memory usage of the model itself, making it a major bottleneck, especially, for memory-constraint devices. In this paper we focus on improving KV cache eviction methods by keeping relevant KV to utilize the memory budget in the best possible way.
Several methods have been proposed to optimize KV cache memory usage, including sparse attention mechanisms that selectively attend to a subset of tokens rather than all previous tokens reducing memory and computational complexity (Zhang et al., 2023; Xiao et al., 2024), efficient attention architectures such as linearized attention (Choromanski et al., 2020), and memory-efficient transformers (Shazeer, 2019) which approximate self-attention to minimize memory consumption. These methods require from-scratch training or model finetuning, however, in this work, we focus on post-training token eviction method which dynamically discard less important tokens from the KV cache to control memory growth (Zhang et al., 2024b). Token eviction methods offer a distinct advantage in memory management by explicitly reducing the KV cache size while preserving critical information. These methods leverage the sparsity of attention — a phenomenon where a small subset of cached keys contribute disproportionately in the attention mechanism — to selectively retain only the most important tokens.
Typically, token importance is determined based on attention scores or, equivalently, attention weights, which measure the alignment between query and key tokens (Zhang et al., 2024b; Oren et al., 2024). However, since the attention output is a linear combination of attention weights and value vectors, evicting tokens simply based on attention scores may lead to suboptimal decisions, as it does not consider the contribution of value vectors from the tokens to be evicted.
In this paper, we propose CAOTE, a post-training cache eviction method that seamlessly integrates eviction (attention) scores with value vectors by optimizing the eviction error, unlike recent methods, as shown in Table 1. CAOTE offers flexibility by being applicable to any post-training eviction method and also includes an efficient counterpart. When combined with recent cache eviction methods (Zhang et al., 2024b; Oren et al., 2024; Li et al., 2024), we observe performance boost for all recent token eviction methods on a variety of downstream tasks: 16 tasks from LongBench, Needle-in-Haystack and Perplexity measurement tasks.
The paper is divided into the following sections: Section 2 discusses background for token eviction. Our main method is in Section 3 which consists of CAOTE. In Section 4, we present experiments and results. Section 5 consists related works and Section 6 consists the conclusion.
| Method | Keys | Values | Minimize eviction error |
| H2O | ✓ | ✗ | ✗ |
| TOVA | ✓ | ✗ | ✗ |
| SnapKV | ✓ | ✗ | ✗ |
| X+CAOTE | ✓ | ✓ | ✓ |
2 Background
Token eviction is a popular methodology (Xiao et al., 2024; Oren et al., 2024; Zhang et al., 2024b) for decoder-only transformer inferences that prevents KV-Cache from growing linearly as token generation continues by preventing less important tokens from being cached. This has dual benefits; first, it limits memory consumption for the KV-cache and second, it reduces the computational complexity of the attention mechanism. Here, we consider the case of processing the input prompt block-wise in resource-restricted environments. In this case, token eviction can save memory and computation not only in the generation phase, but also in the prefill phase, leading to a shorter time-to-first-token which is especially beneficial when the input prompt is extremely long.
With a sequence of hidden states , the transformer block updated the hidden states as follows:
where, is the hidden state of token , denotes the feedforward layer, and denotes the self-attention layer, superscript denotes the layer index.
For brevity, we omit normalization layers and skip connections.
Prompt prefill
Given hidden-states of tokens, the self-attention layer process inputs as follows:
where, and . Here, we omit output layer projection and multi-head extention for brevity.
Block-wise prompt prefill
Instead of processing all tokens at once (resulting in attention matrix: ), we can process tokens in block-size , which also helps in evicting tokens in small blocks instead of larger chunks.
where, the new token hidden states are which are projected to
Generation. In autoregressive generation a single token is generated at each iteration
For resource-constraint hardware, single-inference KV cache prefill for a large number of input tokens may cause out-of-memory error or slow throughput. On the other hand, combining block-wise prefill with token eviction after processing each block of prompt can resolve this issue and improve throughput (Holmes et al., 2024; Agrawal et al., 2023). For a block-size , when tokens are initially processed, the usage of memory and computation power can always be kept within budget constraints by processing the next tokens and evicting the next tokens. In this case, attention matrix has size: .
Recent eviction methods use variants of attention scores from for evicting tokens by using a function (or operator) to map , where is the retention score (or score) for token , the top- tokens are retained based on the score: , where is the budget (maximum tokens allowed per layer). Examples of score functions, for H2O, , and for TOVA, . The process of token eviction follows intuitive steps as shown: computing scores for newly processed tokens, choosing top- tokens, computing attention output using the top- tokens’ hidden-state, we show the steps below:
where the key in bold are correspond to the new tokens’ hidden states being inserted. In above equation we assume that no new query token was evicted for ease of notation. During generation, the flow remains same with .
3 CAOTE: KV Caching through Attention Output-Based Token Eviction
Our method is developed based on two key insights: (i) existing token eviction policies primarily rely on attention scores derived from queries and keys, and (ii) attention output is a linear combination of values. We find that optimizing for eviction error is same as change in attention output due to eviction, which can be computed in closed-form for each token during generation and can be used as the eviction score (CAOTE score).
We first introduce CAOTE in Subsection 3.1 and how to compute eviction error in closed-form. This is followed by a discussion of its meta-property, demonstrating its applicability with other score-based attention methods such as H2O (Zhang et al., 2024b) in Subsection 3.2. Finally, we propose an efficient approximation of CAOTE in Subsection 3.3. The general workflow of CAOTE is illustrated in Fig. 1, highlighting that the modifications to existing token eviction methods are minimal.
3.1 CAOTE Score
The objective of our token eviction is to minimize eviction error: the change in attention output before and after eviction. We formulate eviction error for the generation scenario in which a single new token is inputted and therefore a single token needs to be evicted to maintain the budget . Throughout the paper, we will use eviction error and CAOTE score interchangeably.
Given the attention scores of tokens w.r.t. the last input token and the values: , where is the head dimension. The CAOTE score for token is defined as (we ignore the layer and head dependence for simplicity).
We proof that CAOTE score is same as the eviction error. We define eviction error for token as the mean square error between attention output before and after eviction. Using the same setup as above:
where, is attention output before eviction and is attention output after eviction token .
where, in Eq. (11) is the post-eviction attention score to maintain the sum of the attention score property of sum equal to . In the following we show the relation between the pre and post eviction attention score for token after the eviction of token .
Theorem 3.1.
Given a new input token that exceeds the budget () by . A token needs to be evicted. For any token being evicted, given the retention scores pre-eviction and post-eviction for any token as and respectively, then the following relation holds:
Proof.
Let the last input token has index , then we define
The retention score for token after evicting token is
Theorem 3.2.
Proof.
Using Eq. (19) CAOTE scores (or eviction error) for each token can be computed in parallel as the dependency on only on attention scores and value vectors. Note that this is the firsts formulation seamlessly integrating attention scores and value vectors into a single score. Any norm can be used for computing CAOTE score and based on empirical results we choose -norm.
3.2 CAOTE with general score-based eviction methods
The CAOTE formulation allows the use of arbitrary scoring-based eviction methods to incorporate the values into their scoring mechanism, provided that the scores sum to 1.0. In practice, we can adjust the raw eviction scores without changing their relative order by simple normalizations (affine transformations). Let be the set of retention scores and be the normalizing function. The CAOTE score for general eviction methods is given by:
where, . We further discuss the generalization of CAOTE to well-known token eviction methods in the following.
CAOTE for H2O
We consider H2O (Zhang et al., 2024b), where the scores () are based on the sum of previous attention scores, leading to during generation-phase as proved in Theorem 24 in Section B.1. In this case, simply dividing each token score by the sum of all scores maps the scores to the range and ensures that new scores follow .
For recent methods where all the scores are , simply dividing by sum of all scores suffices. Note that for TOVA (Oren et al., 2024), this summation is by default equal to .
3.3 Fast CAOTE Computation
We also propose a compute-efficient version of CAOTE, FastCAOTE, with negligible performance degradation and reduced computation by an order of , where is the hidden dimension of the model. Here, the pre-eviction attention output () is replaced with mean of values while everything else remains same, that is, CAOTE score for token is:
4 Results
In this section, we demonstrate the efficacy of CAOTE for boosting performance on state-of-the-art token eviction methods on a wide range of downstream benchmarks.
4.1 Experiment Setup
Tasks
We study the impact of CAOTE on different token eviction methods by evaluating on LongBench (Bai et al., 2024), covering single QA, multiple QA, single/multi-document summarization, synthetic, and code generation tasks. We measure long-context perplexity on the Booksum dataset (Kryściński et al., 2021), and lastly, measure recall accuracy on Needle In A Haystack task (Liu et al., 2024; Kamradt, 2023).
Baselines
We compare the performance of CAOTE to various token eviction methods including: H2O (Zhang et al., 2024b), TOVA (Oren et al., 2024), and SnapKV (Li et al., 2024), on LLAMA3 models: Llama 3.2-3B-Instruct and Llama 3.1-8B-Instruct (Dubey et al., 2024), and QWEN2.5 models: Qwen 2.5-3B-Instruct and Qwen 2.5-7B-Instruct (Yang et al., 2024) for all subsequent experiments.
Budgets
We evaluated all methods with various KV cache budget sizes of , , , and , denoted by k, k, k, and k, respectively.
Prompt consumption
Unlike other token eviction methods that assume to prefill prompt at once followed by KV cache eviction, we propose to consume tokens in block-wise manner as described in Section 2 with the block-size of , i.e., at each inference of LLM during the prefill phase, there are new tokens incoming and being added to the cache, and tokens from the cache are evicted once the total number of tokens reaches the cache budget limit.
4.2 LongBench
| Single Doc. QA | Multi Doc. QA | Summarization | Fewshot Learning | Synthetic | Code | |||||||||||||
| Narrative QA | Qasper | MF-en | HotpotQA | 2WikiMQA | Musique | GovReport | QMSum | MultiNews | TREC | TriviaQA | SAMSum | PCount | PR-en | Lcc | RB-P | Avg. | ||
| Llama 3.1-8B | 30.05 | 47.00 | 56.12 | 57.33 | 47.81 | 32.25 | 34.86 | 25.32 | 27.02 | 73.00 | 91.61 | 43.37 | 8.33 | 99.50 | 61.66 | 51.94 | 49.20 | |
| 2k | H2O | 1.74 | 21.15 | 25.33 | 26.11 | 24.15 | 8.78 | 2.17 | 2.70 | 16.78 | 44.00 | 29.36 | 7.62 | 2.25 | 5.88 | 40.15 | 12.14 | 16.89 |
| + CAOTE | 14.32 | 38.34 | 45.97 | 37.77 | 42.51 | 22.06 | 29.57 | 15.11 | 27.02 | 62.00 | 63.60 | 27.34 | 2.00 | 15.50 | 56.99 | 32.87 | 33.31 | |
| + FastCAOTE | 15.15 | 41.27 | 46.6 | 39.91 | 40.02 | 24.55 | 30.05 | 16.19 | 26.95 | 63 | 62.39 | 26.86 | 3.08 | 17.5 | 56.87 | 34.75 | 34.07 | |
| TOVA | 22.57 | 37.26 | 39.43 | 45.74 | 34.48 | 14.77 | 28.87 | 21.17 | 26.95 | 62.50 | 90.73 | 42.74 | 0.00 | 18.00 | 62.68 | 52.48 | 37.52 | |
| + CAOTE | 21.92 | 37.47 | 38.28 | 45.88 | 35.2 | 15 | 29.02 | 21.21 | 27 | 62.5 | 91.34 | 43.22 | 1.5 | 23 | 62.6 | 54.13 | 38.08 | |
| + FastCAOTE | 21.94 | 38.22 | 38.22 | 46.72 | 36.93 | 14.31 | 29.06 | 21.72 | 26.98 | 63 | 91.65 | 43.53 | 1.5 | 22 | 62.44 | 52.88 | 38.19 | |
| SnapKV | 21.81 | 37.22 | 37.19 | 46.10 | 35.42 | 16.53 | 29.83 | 21.05 | 26.77 | 61.00 | 88.84 | 42.56 | 4.03 | 51.50 | 62.37 | 51.45 | 39.60 | |
| + CAOTE | 21.75 | 37.49 | 36.86 | 44.62 | 37.26 | 16.82 | 30.3 | 21.67 | 26.88 | 64 | 90.65 | 42.8 | 2.09 | 53 | 62.5 | 52.09 | 40.05 | |
| + FastCAOTE | 23.26 | 38.54 | 39.16 | 43.2 | 38.27 | 17.54 | 30.28 | 21.97 | 26.76 | 65.5 | 90.91 | 42.71 | 2.84 | 56 | 62.36 | 52.4 | 40.73 | |
| 4k | H2O | 4.07 | 36.16 | 36.00 | 33.52 | 32.87 | 17.78 | 6.66 | 5.95 | 24.09 | 55.00 | 47.65 | 17.41 | 4.00 | 24.50 | 54.85 | 21.43 | 26.37 |
| + CAOTE | 20.28 | 46.08 | 51.45 | 47.38 | 46.05 | 30.89 | 33.39 | 20.8 | 26.93 | 69 | 80.12 | 38.27 | 4.31 | 32 | 59.22 | 40.51 | 40.42 | |
| + FastCAOTE | 24.4 | 44.32 | 48.11 | 48.19 | 43.69 | 21.12 | 31.55 | 22.36 | 26.98 | 65 | 91.18 | 43.11 | 2 | 46.5 | 61.62 | 53.35 | 42.09 | |
| TOVA | 22.68 | 44.55 | 47.87 | 46.76 | 44.54 | 20.56 | 30.95 | 22.13 | 26.96 | 61.50 | 90.56 | 43.27 | 3.00 | 43.50 | 61.62 | 53.40 | 41.49 | |
| + CAOTE | 24.68 | 43.88 | 48.07 | 49.64 | 44.91 | 22.57 | 31.25 | 22.25 | 26.98 | 63 | 91.29 | 43.29 | 2.5 | 46.5 | 61.6 | 53.45 | 42.24 | |
| + FastCAOTE | 24.4 | 44.32 | 48.11 | 48.19 | 43.69 | 21.12 | 31.55 | 22.36 | 26.98 | 65 | 91.18 | 43.11 | 2 | 46.5 | 61.62 | 53.35 | 42.09 | |
| SnapKV | 24.79 | 44.22 | 47.30 | 48.49 | 46.73 | 20.55 | 32.19 | 22.68 | 26.95 | 67.50 | 90.98 | 43.14 | 5.17 | 89.50 | 61.44 | 51.20 | 45.18 | |
| + CAOTE | 24.41 | 43.16 | 47.77 | 50.87 | 44.11 | 21.04 | 32.51 | 22.98 | 26.93 | 69 | 91.31 | 43.18 | 3.33 | 92 | 61.04 | 51.74 | 45.34 | |
| + FastCAOTE | 24.12 | 44.59 | 47.39 | 50.82 | 44.07 | 22.38 | 32.33 | 22.92 | 27.01 | 69 | 91.31 | 43.53 | 4.58 | 94.5 | 61.31 | 52.11 | 45.75 | |
| Llama 3.2-3B | 23.76 | 40.23 | 50.09 | 50.69 | 42.29 | 26.84 | 33.09 | 24.30 | 25.21 | 72.50 | 90.11 | 42.58 | 3.00 | 96.50 | 56.22 | 56.52 | 45.87 | |
| 2k | H2O | 1.63 | 19.96 | 20.20 | 18.02 | 19.56 | 2.88 | 0.78 | 1.55 | 15.97 | 41.00 | 21.97 | 9.83 | 0.50 | 0.50 | 39.71 | 13.91 | 14.25 |
| + CAOTE | 6.38 | 34.36 | 40.6 | 32.52 | 31.08 | 12.69 | 27.36 | 15.04 | 24.6 | 59 | 52.83 | 26.78 | 3.7 | 7.56 | 51.09 | 36.33 | 28.87 | |
| + FastCAOTE | 7.27 | 34.23 | 39.74 | 32.22 | 30.08 | 12.63 | 27.86 | 15.48 | 25.15 | 60.5 | 53.09 | 26.94 | 2.17 | 8.12 | 51.2 | 35.06 | 28.86 | |
| TOVA | 17.14 | 30.14 | 32.44 | 35.96 | 30.05 | 13.08 | 26.15 | 19.70 | 25.04 | 56.50 | 87.81 | 40.48 | 2.50 | 11.50 | 55.51 | 52.36 | 33.52 | |
| + CAOTE | 17.75 | 30.45 | 32.19 | 37.53 | 29.35 | 13.33 | 26.92 | 19.93 | 25.18 | 60.5 | 88.08 | 41.65 | 1.00 | 12.5 | 54.92 | 53.22 | 34.03 | |
| + FastCAOTE | 17.93 | 30.52 | 33.1 | 37.01 | 30.7 | 13.88 | 26.39 | 20.28 | 24.96 | 60.5 | 88.95 | 41.27 | 2.00 | 12.5 | 55.65 | 53.56 | 34.33 | |
| SnapKV | 17.38 | 31.37 | 31.48 | 37.77 | 30.05 | 11.54 | 27.03 | 19.93 | 24.97 | 59.00 | 88.13 | 40.48 | 3.50 | 32.50 | 56.32 | 55.91 | 35.46 | |
| +CAOTE | 19.11 | 33.12 | 31.09 | 38.68 | 32.09 | 12.31 | 27.48 | 20.38 | 25.2 | 64 | 87.7 | 40.78 | 2.5 | 35 | 57.03 | 56.21 | 36.42 | |
| +FastCAOTE | 18.96 | 32.97 | 33.61 | 39.00 | 31.36 | 12.35 | 27.48 | 20.15 | 25.24 | 65 | 87.2 | 40.7 | 4.5 | 36.5 | 56.06 | 57.12 | 36.76 | |
| 4k | H2O | 2.92 | 31.94 | 33.23 | 24.49 | 28.08 | 7.55 | 5.44 | 6.30 | 22.77 | 53.00 | 38.85 | 20.33 | 1.50 | 7.50 | 51.23 | 22.94 | 22.38 |
| +CAOTE | 12.87 | 40.79 | 47.56 | 40.28 | 39.07 | 16.61 | 30.82 | 19.65 | 25.12 | 65.5 | 69.29 | 34.16 | 2.35 | 17 | 55.32 | 45.12 | 35.09 | |
| +FastCAOTE | 11.85 | 40.41 | 47.93 | 40.81 | 38.93 | 17.36 | 31.22 | 19.67 | 25.1 | 65 | 71.25 | 34.89 | 3.5 | 15 | 55.5 | 44.3 | 35.17 | |
| TOVA | 20.52 | 39.53 | 42.47 | 44.12 | 38.42 | 18.22 | 29.36 | 21.36 | 24.96 | 63.50 | 88.98 | 41.50 | 3.00 | 23.50 | 55.72 | 56.66 | 38.24 | |
| +CAOTE | 19.59 | 39.79 | 42.03 | 45.25 | 37.07 | 19.3 | 29.39 | 21.57 | 24.92 | 63 | 89.14 | 41.77 | 3.00 | 29.5 | 55.68 | 56.19 | 38.57 | |
| +FastCAOTE | 19.77 | 39.23 | 43.13 | 45.28 | 37.04 | 18.82 | 29.25 | 21.94 | 24.96 | 63 | 88.64 | 41.92 | 3.5 | 29 | 55.68 | 56.41 | 38.6 | |
| SnapKV | 19.85 | 39.22 | 39.86 | 46.70 | 37.98 | 16.64 | 29.79 | 21.21 | 25.01 | 65.50 | 89.35 | 40.95 | 2.50 | 62.50 | 55.74 | 56.88 | 40.60 | |
| +CAOTE | 20.1 | 39.74 | 41.01 | 45.64 | 38.26 | 18.64 | 30.07 | 21.53 | 24.98 | 67.5 | 89.08 | 41.78 | 3.00 | 67 | 55.73 | 56.71 | 41.30 | |
| +FastCAOTE | 19.68 | 39.24 | 41.03 | 44.52 | 39.09 | 18.62 | 30.15 | 21.72 | 24.97 | 67 | 88.86 | 41.24 | 3.00 | 71 | 55.67 | 56.64 | 41.40 | |
| Single Doc. QA | Multi Doc. QA | Summarization | Fewshot Learning | Synthetic | Code | |||||||||||||
| Narrative QA | Qasper | MF-en | HotpotQA | 2WikiMQA | Musique | GovReport | QMSum | MultiNews | TREC | TriviaQA | SAMSum | PCount | PR-en | Lcc | RB-P | Avg. | ||
| Qwen 2.5-7B | 15.75 | 16.94 | 32.38 | 11.89 | 11.88 | 7.95 | 34.33 | 19.91 | 22.67 | 65.5 | 87.05 | 44.75 | 4.22 | 93.08 | 57.74 | 61.84 | 36.74 | |
| 2k | H2O | 2.39 | 7.29 | 12.42 | 8.55 | 11.06 | 2.73 | 3.62 | 6.6 | 15.69 | 42.5 | 28.21 | 10.63 | 0.65 | 0 | 35.1 | 18.77 | 12.89 |
| + CAOTE | 4.55 | 14.3 | 27.58 | 11.33 | 13.55 | 7.76 | 26.65 | 15.62 | 22.93 | 57 | 49.78 | 27.74 | 1.54 | 11.08 | 51.45 | 32.7 | 23.47 | |
| + FastCAOTE | 4.8 | 12.79 | 28.72 | 12.94 | 13.25 | 7.53 | 27.06 | 14.46 | 22.84 | 59 | 48.23 | 26.4 | 2.53 | 11.54 | 52.85 | 32.93 | 23.62 | |
| TOVA | 8.49 | 14.01 | 21.04 | 14 | 11.51 | 5.09 | 27.43 | 17.84 | 22.83 | 56.5 | 79.56 | 40.55 | 2.43 | 9.29 | 55.99 | 56.15 | 27.67 | |
| + CAOTE | 10.46 | 14.82 | 25.06 | 14.62 | 11.73 | 6.01 | 27.66 | 18.02 | 22.78 | 57.5 | 79.39 | 40.87 | 2.5 | 11.25 | 56.22 | 56.51 | 28.46 | |
| + FastCAOTE | 10.08 | 13.58 | 25.28 | 14.44 | 12.14 | 5.24 | 27.34 | 18.31 | 23.11 | 55.5 | 78.51 | 41.67 | 2.7 | 10.54 | 56.56 | 58.05 | 28.32 | |
| SnapKV | 11.6 | 12.45 | 23.66 | 12.38 | 10.64 | 7.03 | 27.57 | 18.27 | 22.85 | 58 | 81.78 | 41.13 | 3.76 | 19.42 | 55.83 | 56.53 | 28.93 | |
| + CAOTE | 14.02 | 12.23 | 24.55 | 16.45 | 10.35 | 8.59 | 27.77 | 18.91 | 22.87 | 56 | 80.58 | 40.43 | 2.38 | 21.52 | 55.17 | 56.03 | 29.24 | |
| + FastCAOTE | 14.26 | 14.11 | 24.11 | 15.31 | 11.35 | 7.88 | 27.95 | 18.86 | 22.74 | 56.5 | 80.92 | 41.49 | 3.8 | 22.42 | 55.89 | 57.43 | 29.69 | |
| 4k | H2O | 1.99 | 11.92 | 19.88 | 10.24 | 10.12 | 4.73 | 9.08 | 10.14 | 20.85 | 51.00 | 37.37 | 20.57 | 3.16 | 6.43 | 52.14 | 29.09 | 18.67 |
| + CAOTE | 4.78 | 18.06 | 32.49 | 16.23 | 17.28 | 9.57 | 29.81 | 18.04 | 22.86 | 59.5 | 63.05 | 36.91 | 2.7 | 28.25 | 55.13 | 42.42 | 28.57 | |
| + FastCAOTE | 5.69 | 16.99 | 32.62 | 18.22 | 16.58 | 10.48 | 30.3 | 17.71 | 22.88 | 59.5 | 62.95 | 36.29 | 2.1 | 27.65 | 56.3 | 40.65 | 28.56 | |
| TOVA | 12.83 | 17.03 | 27.01 | 16.8 | 13.37 | 8.05 | 29.21 | 19.05 | 22.73 | 58.5 | 82.67 | 42.71 | 1.67 | 15 | 56.69 | 56.59 | 29.99 | |
| + CAOTE | 12.97 | 14.99 | 27.53 | 17.94 | 12.93 | 9.21 | 29.76 | 19.7 | 22.92 | 58 | 82.03 | 43.14 | 2.15 | 17.25 | 57.32 | 59.37 | 30.98 | |
| + FastCAOTE | 14.52 | 16.71 | 26.97 | 18.73 | 13.84 | 9.59 | 29.47 | 19.45 | 22.87 | 59.5 | 82.96 | 42.42 | 2.6 | 20.33 | 57.22 | 58.42 | 30.98 | |
| SnapKV | 14.35 | 13.45 | 28.28 | 16.33 | 11.74 | 8.12 | 29.71 | 19.18 | 22.82 | 57 | 83.8 | 43.27 | 2.41 | 39.83 | 58.12 | 58.67 | 31.69 | |
| + CAOTE | 15.07 | 14.34 | 28.7 | 16.7 | 12.89 | 10.54 | 30.03 | 19.58 | 22.73 | 59.5 | 83.12 | 42.56 | 3.17 | 55.92 | 57.34 | 58.85 | 33.19 | |
| + FastCAOTE | 17.12 | 14.69 | 27.6 | 17.52 | 13.69 | 9.96 | 30.24 | 20.02 | 22.88 | 58.5 | 81.13 | 42.31 | 4.06 | 53.33 | 57.51 | 58.77 | 33.08 | |
| Qwen 2.5-3B | 18.08 | 22.49 | 39.72 | 27.86 | 20.45 | 18.93 | 32.8 | 23.74 | 24.89 | 67.5 | 85.05 | 43.88 | 5 | 40.97 | 51.91 | 47.53 | 35.68 | |
| 2k | H2O | 1.8 | 9.18 | 11.62 | 8.54 | 7.31 | 2.77 | 5.93 | 6.99 | 16.89 | 38 | 21.87 | 7.69 | 1 | 3 | 37.36 | 22.9 | 12.68 |
| + CAOTE | 6.9 | 22.71 | 28.09 | 15.23 | 18.19 | 4.95 | 29.53 | 17.68 | 24.74 | 52.5 | 45.81 | 26.95 | 1.92 | 6.16 | 45.81 | 36.48 | 23.98 | |
| + FastCAOTE | 7.03 | 22.37 | 28.88 | 15.34 | 16.95 | 5.19 | 29.13 | 18.06 | 25.03 | 54.5 | 46.35 | 25.35 | 2.23 | 7.22 | 45.7 | 36.59 | 24.12 | |
| TOVA | 11.69 | 14.94 | 25.33 | 17.29 | 12.58 | 5.91 | 26.67 | 21.49 | 24.78 | 51.5 | 68.8 | 41.79 | 0.23 | 6 | 49.79 | 48.6 | 26.71 | |
| + CAOTE | 11.17 | 15.23 | 27.42 | 18.94 | 13.1 | 6.94 | 27.01 | 21.62 | 24.86 | 57.5 | 68.38 | 42.11 | 0.82 | 4.88 | 49.36 | 48.13 | 27.34 | |
| + FastCAOTE | 11.04 | 15.36 | 27.72 | 19.8 | 13.65 | 6.37 | 27.17 | 22.08 | 24.64 | 57 | 69.13 | 42.48 | 0.77 | 5.25 | 48.36 | 48.58 | 27.46 | |
| SnapKV | 11.7 | 13.91 | 24.28 | 14.8 | 10.89 | 7.42 | 27.4 | 21.63 | 24.64 | 54.5 | 75.35 | 42.72 | 2.5 | 18.33 | 49.65 | 50.59 | 28.14 | |
| +CAOTE | 12.69 | 14.88 | 26.16 | 13.93 | 12.21 | 7.07 | 27.48 | 20.99 | 24.75 | 61 | 75.58 | 42.08 | 4 | 21.29 | 49.94 | 52.38 | 29.15 | |
| +FastCAOTE | 12.03 | 14.56 | 24.82 | 14.66 | 10.83 | 7.89 | 27.51 | 20.98 | 24.67 | 62.5 | 75.51 | 41.53 | 2 | 17 | 49.02 | 50.83 | 28.52 | |
| 4k | H2O | 2.82 | 17.34 | 23.27 | 10.18 | 10.47 | 3.03 | 11.06 | 10.73 | 22.93 | 50.75 | 34.93 | 18.03 | 4.35 | 7.32 | 47.74 | 29.42 | 19.02 |
| +CAOTE | 7.63 | 24.16 | 35.29 | 20.17 | 17.67 | 12.61 | 31.14 | 19.04 | 25.01 | 62.5 | 64.84 | 34.19 | 4.25 | 18.37 | 49.79 | 41.45 | 29.26 | |
| +FastCAOTE | 8.58 | 23.45 | 33.14 | 21.72 | 16.11 | 12.26 | 31.11 | 19.76 | 25.04 | 62 | 65.01 | 35.15 | 4.6 | 17.88 | 50.05 | 40.03 | 29.12 | |
| TOVA | 12.19 | 18.31 | 32.56 | 20.58 | 13.8 | 7.74 | 28.82 | 22.27 | 24.98 | 59 | 80.66 | 43.05 | 1.11 | 9.56 | 49.93 | 46.74 | 29.46 | |
| +CAOTE | 13.16 | 18.67 | 30.74 | 19.33 | 15.7 | 7.32 | 28.93 | 22.14 | 24.91 | 59.5 | 78.54 | 43.57 | 1.55 | 8.25 | 49.4 | 47.31 | 29.31 | |
| +FastCAOTE | 12.2 | 18.55 | 32.29 | 19.3 | 15.13 | 7.23 | 29.12 | 22.44 | 24.97 | 60 | 78.8 | 43.12 | 1.5 | 10.25 | 49.6 | 47.74 | 29.52 | |
| SnapKV | 12.98 | 2.21 | 31.77 | 18.33 | 14.41 | 10.83 | 29.14 | 22.38 | 24.89 | 61 | 84.17 | 42.63 | 3.75 | 25.42 | 50.22 | 48.77 | 30.18 | |
| +CAOTE | 13.65 | 20.35 | 32.62 | 19.36 | 15.27 | 11.42 | 29.47 | 22.44 | 24.78 | 64 | 82.6 | 43 | 4.25 | 24.46 | 50.37 | 49.28 | 31.71 | |
| +FastCAOTE | 13.46 | 19.92 | 32.53 | 20.44 | 13.64 | 8.44 | 29.56 | 22.14 | 24.93 | 64.5 | 82.73 | 43.26 | 2 | 24.58 | 50.14 | 50.08 | 31.40 | |
We present the accuracy of Llama 3.1-8B-Instruct, Llama 3.2-3B-Instruct and, Qwen 2.5-3B-Instruct, Qwen 2.5-7B-Instruct using baseline eviction methods with budget of k, k, both with and without CAOTE in Table 2 and Table 3. We observe that the best performance is given by SnapKV-FastCAOTE for the Llama3 models, while for Qwen 2.5 models SnapKV-CAOTE performs the best. H2O shows improvement with CAOTE, while TOVA, SnapKV also show overall improvements, making their average accuracy closer to dense accuracy. Additional results for the k, k budget are shown in Table 7, Table 8 for Llama3 and Qwen 2.5 respectively, in Section C.1, which follow a trend similar to the k, k budgets.
4.3 Perplexity
| Budget | H2O | TOVA | SnapKV | ||||||
| +CAOTE | +FastCAOTE | +CAOTE | +FastCAOTE | +CAOTE | +FastCAOTE | ||||
| Llama 3.1-8B-Instruct | |||||||||
| 2k | 2.007 | 1.884 | 1.891 | -0.046 | -0.088 | -0.085 | -0.019 | -0.097 | -0.098 |
| 4k | 1.284 | 1.079 | 1.061 | -0.047 | -0.060 | -0.058 | -0.0483 | -0.080 | -0.079 |
| 6k | 0.843 | 0.716 | 0.703 | -0.035 | -0.0366 | -0.085 | -0.036 | -0.043 | -0.045 |
| Llama 3.2-3B-Instruct | |||||||||
| 2k | 3.814 | 3.561 | 3.563 | 0.493 | 0.442 | 0.432 | 0.555 | 0.451 | 0.435 |
| 4k | 2.460 | 2.142 | 2.128 | 0.175 | 0.150 | 0.144 | 0.223 | 0.152 | 0.144 |
| 6k | 1.369 | 1.219 | 1.187 | 0.065 | 0.057 | 0.057 | 0.076 | 0.056 | 0.044 |
| 8k | 0.589 | 0.462 | 0.448 | 0.023 | 0.012 | 0.011 | 0.020 | 0.007 | 0.005 |
We use the Booksum dataset (Kryściński et al., 2021) to measure generation perplexity of different eviction methods for various budgets. In Table 4, we show perplexity gap between a model using a given eviction strategy and that of the model without token eviction with cache budgets of k, k and k. We observe that when CAOTE is applied to existing eviction methods, the perplexity either improves or surpasses the perplexity of the baseline model. TOVA-FastCAOTE, SnapKV-CAOTE, and SnapKV-FastCAOTE perform best for k, k, k budgets, respectively, for Llama 3.1-8B-Instruct; for Llama 3.2-3B-Instruct, TOVA-FastCAOTE performs best with k and k budgets and SnapKV-FastCAOTE beats other methods using k and k. Perplexity results for Qwen 2.5 models are shown in Table 9 in Section C.2.
4.4 Needle In A HayStack
| Budget | H2O | TOVA | SnapKV | ||||||
| +CAOTE | +FastCAOTE | +CAOTE | +FastCAOTE | +CAOTE | +FastCAOTE | ||||
| Llama 3.1-8B-Instruct | |||||||||
| 2k | 0.174 | 0.270 | 0.264 | 0.196 | 0.204 | 0.202 | 0.214 | 0.226 | 0.242 |
| 4k | 0.330 | 0.538 | 0.568 | 0.286 | 0.298 | 0.292 | 0.360 | 0.392 | 0.420 |
| 6k | 0.544 | 0.698 | 0.676 | 0.370 | 0.402 | 0.396 | 0.490 | 0.550 | 0.580 |
| Llama 3.2-3B-Instruct | |||||||||
| 2k | 0.104 | 0.160 | 0.172 | 0.172 | 0.150 | 0.166 | 0.154 | 0.172 | 0.168 |
| 4k | 0.198 | 0.262 | 0.294 | 0.220 | 0.232 | 0.232 | 0.226 | 0.222 | 0.232 |
| 6k | 0.258 | 0.308 | 0.322 | 0.258 | 0.278 | 0.270 | 0.272 | 0.264 | 0.312 |
| 8k | 0.324 | 0.414 | 0.404 | 0.338 | 0.364 | 0.344 | 0.342 | 0.336 | 0.366 |
Lastly, we run extensive experiments on Needle-In-A-Haystack benchmark (Liu et al., 2024; Kamradt, 2023) and show quantitative results in Table 5 and visualizations for k budget Llama 3.1-8B model in Figure 2. We observe in Table 5 that H2O-FastCAOTE performs best for all budgets with Llama 3.2-3B. When using a budget k with Llama 3.1-8B, CAOTE boosted H2O outperforms TOVA, SnapKV as well. H2O-CAOTE performs best for Llama 3.1-8B with budget = {k, k} and H2O-FastCAOTE performs best fork budget for Llama3.1-8b. The gains in precision are especially high for the k budget for the Llama 3.1-8B model, with an increase of up to . We can see in Figure 2 that CAOTE improves the state-of-the-art eviction method and is able to predict beyond their budget constraints. Results for Qwen 2.5 models are shown in Table 10 in Section C.3.
4.5 Ablations
Difference in Attention Outputs
In Figure 3, we check the average change in attention score in different layers for various eviction methods to provide insight into the performance gain realized by CAOTE. We measure the difference between the dense attention-output and eviction method based attention output, both with and without CAOTE. We repeat this for each layer and average the result to verify that CAOTE results in smaller distance to dense attention-output. Figure 3 shows the normalized mean squared error between the dense attention output and eviction-based attention output for H2O, TOVA and SnapKV, both with and without CAOTE in dashed-lines and solid-lines respectively. We observe that the CAOTE variant always produces a smaller gap from dense attention output (no eviction) compared to baselines, supporting our claim that CAOTE evicts the token which minimizes deviation from true attention output. Furthermore, we show that discrepancy in attention output due to eviction (eviction error) causes discrepancy in logits in B.2; to emphasize that optimizing for eviction error results in less discrepancy in logits and, therefore, less deviation on downstream task accuracy.
Different Prompt Block Sizes
| Task | H2O | H2O-CAOTE | TOVA | TOVA-CAOTE | SnapKV | SnapKV-CAOTE | ||||||||||||
| block-size | 32 | 128 | 32 | 128 | 32 | 128 | 32 | 128 | 32 | 128 | 32 | 128 | ||||||
| HotPotQA∗ | 20.43 | 24.49 | 35.71 | 32.37 | 40.28 | 46.37 | 44.36 | 45.55 | 52.93 | 44.72 | 45.25 | 52.94 | 45.99 | 46.7 | 53.67 | 44.66 | 45.64 | 52.67 |
| Qasper | 30.8 | 31.94 | 32.58 | 36.74 | 40.79 | 38.93 | 38.65 | 38.83 | 40.45 | 39.39 | 39.79 | 39.90 | 38.71 | 39.22 | 39.80 | 40.18 | 39.74 | 39.69 |
| GovReport | 5.52 | 5.44 | 7.07 | 30.10 | 30.82 | 31.48 | 29.32 | 29.41 | 31.25 | 29.24 | 29.39 | 31.15 | 29.62 | 29.79 | 31.82 | 30.15 | 30.07 | 32.42 |
| Trec | 46.50 | 53.00 | 56.00 | 59.00 | 65.5 | 72.00 | 62.00 | 64.00 | 71.50 | 64.00 | 63.00 | 71.50 | 64.00 | 65.50 | 71.50 | 67.50 | 67.50 | 71.50 |
| LCC | 51.59 | 51.23 | 51.51 | 54.06 | 55.32 | 54.51 | 56.23 | 55.64 | 56.34 | 56.18 | 55.68 | 56.39 | 56.24 | 55.74 | 56.56 | 56.27 | 55.73 | 56.59 |
Prompt block size dictates how many tokens will be evicted at once. A moderately large prompt block size results in faster prompt processing but evicts more tokens each time, while a smaller prompt block size results in slower prefilling. To evaluate the impact of prompt block size, we evaluated a subset of LongBench tasks with block prompt sizes of , 128 and , i.e., process the entire prompt as a single large block and display the results in Table 6 for Llama3.2 with budget=k. Based on the CAOTE formulation (single token eviction), the smaller the block-size the better CAOTE would perform as seen for block-size= in Table 6, however, even for large prompt block size, CAOTE performs competitive to baseline if not better. Note that the infinite prompt block size initially violates the memory/budget constraint, and additionally, we observe out-of-memory errors for several samples for the HotPotQA and GovReport tasks when using infinite block size. This problem is not present with block-size=.
5 Related Work
Sparse and Efficient Attention
Sparse or efficient attention based methods result in mitigating the computation load and saving memory consumption by using efficient linear attentions (Katharopoulos et al., 2020). Additionally, there are KV compression methods which don’t evict any tokens as post eviction the token is not retrievable, (Dong et al., 2024) proposes to keep important tokens based on attention score in cache while combining the evicted tokens via linear attention into single embedding. Landmark attention injects learnable special tokens between chunks of tokens and access past tokens in chunks instead of individually. Lastly, there are better architectures based with constant KV memory which outperform linear attention attentions (Mohtashami & Jaggi, 2023). However, all these methods require either from-scratch training or fine-tuning.
KV Cache Eviction
At the extreme end of efficient KV cache management, token eviction methods have been extensively studied. Leveraging the sparsity of attention in LLMs (Xiao et al., 2024; Sheng et al., 2023; Chen et al., 2021), these methods determine the importance of KV pairs using (learned) rules and retain the pairs with high scores in the cache to approximate the attention output. StreamingLLM (Xiao et al., 2024) observes an attention sink phenomenon, which states that the first few tokens tend to receive the majority of attention weights. To exploit this, it proposes SinkAttention, which prioritizes keeping the initial tokens in the cache while doing sliding window-based attention. Other methods, such as H2O (Zhang et al., 2024b), TOVA (Oren et al., 2024), SnapKV (Li et al., 2024), and RoCO (Ren & Zhu, 2024), retain tokens with high attention scores in the KV cache with various algorithmic modifications. These include preserving the first or last tokens in addition to those with high attention scores or applying smoothing techniques to the attention scores. While these token eviction methods primarily rely on attention scores to assess token importance, CAOTE introduces an orthogonal scoring metric that estimates the impact of values on approximating attention outputs. This metric can complement existing token importance scoring approaches, enhancing other eviction methods.
Furthermore, other lines of work enable layer-wise budget optimization (Feng et al., 2024) by considering the scores of all heads jointly, and selecting top-K nodes, while others (Ge et al., 2023) consider managing memory by keeping/discarding based on token characteristics with baseline token eviction (H2O). Our proposed method, CAOTE is highly flexible, and can be integrated with both the mentioned methods to achieve additional boost on performance.
6 Conclusion
We propose a post-training KV cache eviction method that can be seamlessly integrated with any existing eviction strategies. Our approach, CAOTE, introduces an optimization objective aimed at minimizing the alteration in attention output when evicting a token. This objective ensures the incorporation of both attention scores and value vectors in the eviction decision process. Our formulation allows for the parallel computation of the CAOTE score for all tokens. Additionally, we present an efficient variant, FastCAOTE. Through extensive evaluations across various downstream tasks, we demonstrate that eviction methods equipped with CAOTE consistently deliver superior performance.
Impact Statement
This paper presents work whose goal is to advance the field of Machine Learning. There are many potential societal consequences of our work, none of which we feel must be specifically highlighted here.
Appendix A Limitation
CAOTE is a myopic (greedy) strategy and its scoring framework based on the assumption of evicting token per iteration. This assumption breaks during prefilling stage, however, taking into account change in attention output due to multi-token eviction is non-trivial. Fortunately, we observe that even assuming multi-token eviction independently (without considering the effect of other tokens being evicted), CAOTE is still able to give boost in performance for all tasks. Due to this reason CAOTE’s performance might further improve with smaller prompt filling block-size.
Appendix B Additional Proofs
B.1 H2O scores
Theorem B.1.
Given H2O scores for tokens as , the summation of all , is greater than
Proof.
Assuming that only tokens are present and are propagated through the model at the same time. The causal attention mask , will have all entries on upper triangle excluding diagonal is . The first token will attend to itself have attention score as
H2O score for token is defined as the sum of attention score to token by all future tokens:
In general for a token , the H2O score is defined as
B.2 Relation of CAOTE score (eviction error) to logits
We show that evicting token based on CAOTE score can lead to smaller discrepancy in final logits which affect the downstream performance. As CAOTE score is the eviction error during generation phase, we instead show the relation between eviction error and logits. We start by showing for a single attention layer (single head) based network, its extension to multiple heads and, finally a transformer layer (self-attention and feed-forward-network). For simplicity we ignore layer-norms. Some definitions which will used are given below. We assume a budget of , with current token sequence having tokens (superscript denotes layer):
The attention output for a sequence of tokens is (for layer )
The logits with and without eviction for token are defined as respectively.
(Case 1) Single self-attention layer with single head: The logits for dense baseline is:
where are the LM-head, output projection, and value projection respectively. is the output after the residual connection.
The logits with eviction will have a perturbation due to error in attention output (CAOTE score or eviction error), and is given as:
where the logit error is , from (9).
(Case 2) Multiple attention heads: this is trivial and can be achieved by replacing , where super-script denotes head number.
(Case 3) Single self-attention and feedforward-network (FFN): we still assume single head without layer-norms. The dense logit is given as
where, for simplicity we assume feedforward network to subsumed within .
The perturbed logit due to eviction is given as:
where, the logit error . Thus, the above analysis shows that error in attention output can cause discrepancy in logit space which can affect performance on downstream tasks.
Note that for multiple layers, each layer would have its own eviction error which will keep compounding; however, this computing the exact compounded error is non-trivial due to the presence of output layer-norms.
Appendix C Additional Results
C.1 LongBench
| Single Doc. QA | Multi Doc. QA | Summarization | Fewshot Learning | Synthetic | Code | |||||||||||||
| Narrative QA | Qasper | MF-en | HotpotQA | 2WikiMQA | Musique | GovReport | QMSum | MultiNews | TREC | TriviaQA | SAMSum | PCount | PR-en | Lcc | RB-P | Avg. | ||
| Llama 3.1-8B | 30.05 | 47.00 | 56.12 | 57.33 | 47.81 | 32.25 | 34.86 | 25.32 | 27.02 | 73.00 | 91.61 | 43.37 | 8.33 | 99.50 | 61.66 | 51.94 | 49.20 | |
| 6k | Sink | 25.41 | 47.40 | 44.13 | 47.39 | 45.73 | 21.90 | 32.53 | 22.19 | 26.87 | 72.00 | 91.25 | 43.41 | 3.08 | 52.50 | 62.22 | 56.24 | 43.39 |
| H2O | 8.52 | 43.31 | 44.80 | 40.03 | 42.46 | 21.68 | 11.85 | 8.78 | 26.03 | 62.00 | 56.39 | 25.72 | 5.75 | 45.50 | 58.62 | 29.53 | 33.19 | |
| + CAOTE | 25.77 | 46.45 | 54.99 | 50.57 | 47.7 | 31.93 | 33.99 | 22.86 | 27.01 | 71 | 87.05 | 41.29 | 6.67 | 54 | 60.48 | 43.23 | 44.06 | |
| + FastCAOTE | 27.02 | 46.46 | 55.4 | 51.32 | 47.4 | 32.89 | 34.08 | 23.69 | 27.03 | 71 | 86.25 | 42.14 | 9 | 48.5 | 60.49 | 42.94 | 44.10 | |
| TOVA | 24.59 | 45.93 | 53.92 | 55.09 | 47.43 | 25.07 | 32.33 | 24.10 | 27.00 | 68.50 | 90.81 | 43.89 | 4.25 | 67.00 | 61.50 | 52.39 | 45.24 | |
| + CAOTE | 24.23 | 45.88 | 53.5 | 52.96 | 49.59 | 27.02 | 32.62 | 23.86 | 27.08 | 70 | 90.98 | 43.45 | 3 | 74.5 | 61.46 | 51.76 | 45.74 | |
| + FastCAOTE | 24.17 | 46.07 | 53.8 | 53.53 | 48.11 | 26.49 | 32.64 | 23.88 | 27.01 | 70 | 90.81 | 43.53 | 3 | 73 | 61.49 | 51.43 | 45.56 | |
| SnapKV | 24.10 | 45.57 | 50.44 | 53.12 | 48.41 | 24.27 | 33.43 | 23.53 | 27.03 | 71.50 | 92.28 | 43.58 | 5.25 | 98.00 | 61.32 | 52.16 | 47.12 | |
| + CAOTE | 25.97 | 46.09 | 51.54 | 55.19 | 47.41 | 26.48 | 33.32 | 24 | 27.05 | 71.5 | 91.11 | 43.55 | 6.83 | 99.5 | 61.11 | 51.45 | 47.63 | |
| + FastCAOTE | 24.77 | 46.06 | 52.18 | 56.72 | 47.01 | 26.24 | 33.41 | 23.8 | 26.99 | 73 | 91.31 | 43.6 | 5.92 | 99.5 | 61.5 | 51.37 | 47.71 | |
| 8k | Sink | 23.53 | 46.63 | 48.68 | 49.61 | 47.16 | 21.14 | 33.10 | 23.20 | 26.92 | 72.00 | 91.29 | 43.79 | 3.25 | 66.00 | 62.18 | 56.43 | 44.68 |
| H2O | 13.85 | 44.94 | 47.81 | 43.64 | 44.90 | 23.65 | 18.78 | 11.35 | 26.49 | 69.50 | 69.05 | 33.41 | 5.25 | 62.50 | 59.74 | 36.26 | 38.20 | |
| + CAOTE | 27.74 | 46.67 | 54.97 | 52.71 | 48.28 | 33.66 | 34.51 | 24.73 | 26.99 | 73 | 86.8 | 42.86 | 5 | 66.5 | 61.06 | 48.29 | 45.86 | |
| + FastCAOTE | 28.8 | 47.08 | 54.67 | 52.95 | 47.13 | 34.36 | 34.21 | 24.53 | 27.04 | 72 | 87.87 | 43.03 | 5.5 | 69.5 | 61.06 | 49.69 | 46.21 | |
| TOVA | 24.86 | 46.78 | 54.83 | 54.52 | 49.00 | 26.40 | 33.44 | 24.76 | 27.00 | 71.00 | 91.11 | 43.29 | 6.25 | 87.00 | 61.49 | 51.79 | 47.09 | |
| + CAOTE | 25.65 | 46.88 | 54.5 | 55.42 | 48.73 | 26.54 | 33.47 | 24.8 | 27.02 | 72 | 91.11 | 43.26 | 6.25 | 87 | 61.36 | 51.3 | 47.21 | |
| + FastCAOTE | 25.25 | 46.75 | 54.76 | 56.29 | 48.94 | 26.25 | 33.37 | 24.81 | 27.01 | 71.5 | 91.11 | 43.24 | 5.25 | 89 | 61.32 | 52.1 | 48.58 | |
| SnapKV | 25.15 | 46.55 | 53.39 | 56.00 | 48.75 | 27.82 | 33.67 | 24.85 | 27.01 | 72.50 | 91.78 | 43.54 | 5.08 | 100.00 | 61.48 | 51.41 | 48.06 | |
| + CAOTE | 27.06 | 46.42 | 53.76 | 56.51 | 47.79 | 28.13 | 33.87 | 24.93 | 27.02 | 73 | 91.38 | 43.29 | 6.75 | 99.5 | 61.49 | 51.99 | 48.31 | |
| + FastCAOTE | 26.91 | 46.59 | 53.47 | 56.63 | 48.54 | 29.27 | 33.91 | 24.86 | 27.01 | 73 | 91.38 | 43.49 | 6.75 | 100 | 61.49 | 51.78 | 48.44 | |
| Llama 3.2-3B | 23.76 | 40.23 | 50.09 | 50.69 | 42.29 | 26.84 | 33.09 | 24.30 | 25.21 | 72.50 | 90.11 | 42.58 | 3.00 | 96.50 | 56.22 | 56.52 | 45.87 | |
| 6k | Sink | 19.33 | 40.29 | 37.95 | 46.48 | 40.29 | 15.31 | 30.43 | 21.35 | 25.14 | 71.50 | 88.93 | 42.04 | 3.50 | 47.00 | 56.55 | 54.11 | 40.01 |
| H2O | 4.62 | 38.81 | 39.06 | 34.66 | 35.52 | 15.21 | 10.51 | 10.01 | 24.25 | 61.50 | 53.23 | 27.37 | 0.50 | 13.00 | 54.55 | 32.29 | 28.44 | |
| +CAOTE | 16.14 | 41.68 | 49.36 | 46.7 | 43.36 | 22.75 | 32.07 | 21 | 25.07 | 69 | 80.02 | 39.33 | 1.5 | 26 | 55.82 | 49.05 | 38.68 | |
| +FastCAOTE | 16.31 | 41.94 | 49.17 | 45.64 | 41.83 | 21.68 | 32.07 | 20.73 | 25.02 | 68.5 | 80.34 | 39.88 | 3.5 | 24 | 55.83 | 48.7 | 38.45 | |
| TOVA | 20.22 | 39.78 | 45.86 | 49.08 | 41.54 | 20.43 | 30.50 | 22.17 | 25.11 | 66.50 | 89.00 | 42.50 | 4.00 | 46.50 | 55.57 | 57.53 | 41.02 | |
| +CAOTE | 21.17 | 39.69 | 47.21 | 48.82 | 41.7 | 20.59 | 30.72 | 22.36 | 25.1 | 68 | 89 | 42.38 | 3.5 | 52.5 | 55.6 | 57.09 | 41.59 | |
| +FastCAOTE | 21.48 | 39.66 | 47.02 | 47.56 | 41.95 | 19.91 | 30.8 | 21.98 | 25.17 | 67.5 | 89.5 | 42.06 | 4 | 53 | 55.6 | 57.39 | 41.54 | |
| SnapKV | 20.83 | 39.65 | 44.48 | 49.30 | 40.18 | 20.28 | 31.27 | 22.73 | 25.09 | 69.00 | 89.95 | 41.47 | 4.00 | 85.00 | 55.69 | 57.82 | 43.55 | |
| +CAOTE | 20.23 | 39.65 | 44.91 | 50.16 | 40.58 | 21.32 | 31.23 | 22.51 | 25.13 | 69 | 90 | 41.83 | 5 | 89.5 | 55.84 | 57.24 | 44.01 | |
| +FastCAOTE | 20.09 | 40.02 | 44.58 | 48.57 | 42.12 | 22.51 | 31.25 | 22.89 | 25.15 | 71 | 90 | 41.83 | 4 | 89.5 | 55.83 | 57.25 | 44.16 | |
| 8k | Sink | 20.15 | 40.02 | 41.94 | 48.15 | 42.24 | 16.01 | 31.64 | 22.10 | 25.20 | 73.00 | 89.26 | 42.37 | 3.50 | 62.50 | 56.86 | 56.63 | 41.97 |
| H2O | 9.65 | 39.66 | 43.20 | 38.09 | 40.41 | 21.46 | 17.80 | 13.28 | 24.67 | 70.00 | 64.30 | 32.19 | 2.00 | 24.50 | 55.00 | 39.09 | 33.46 | |
| +CAOTE | 20.07 | 40.73 | 47.76 | 47.25 | 42.88 | 23.19 | 32.41 | 22.01 | 25.15 | 71 | 83.58 | 40.8 | 3 | 43.5 | 55.45 | 53.35 | 40.76 | |
| +FastCAOTE | 20.81 | 40.54 | 48.1 | 47.35 | 43.4 | 25.13 | 32.73 | 22.31 | 25.18 | 71.5 | 84.91 | 40.6 | 4 | 45 | 55.84 | 52.89 | 41.27 | |
| TOVA | 21.08 | 40.67 | 49.07 | 48.69 | 41.93 | 23.05 | 31.64 | 22.85 | 25.21 | 69.00 | 89.25 | 42.19 | 2.50 | 71.00 | 55.77 | 57.47 | 43.21 | |
| +CAOTE | 21.97 | 40.66 | 49.37 | 50.1 | 41.29 | 24.05 | 31.65 | 22.85 | 25.16 | 69.5 | 89.5 | 42 | 3 | 78.5 | 55.82 | 57.16 | 43.91 | |
| +FastCAOTE | 22.73 | 40.51 | 49.36 | 50.18 | 42.26 | 24.45 | 31.68 | 23.09 | 25.16 | 69.5 | 89.5 | 42.28 | 4.5 | 80 | 55.79 | 57.16 | 44.26 | |
| SnapKV | 20.49 | 40.80 | 48.16 | 48.78 | 41.65 | 24.79 | 31.81 | 23.46 | 25.17 | 70.00 | 90.17 | 41.99 | 5.00 | 94.00 | 55.77 | 57.29 | 44.96 | |
| +CAOTE | 19.71 | 40.7 | 48.05 | 49.03 | 41.27 | 22.95 | 31.95 | 23.1 | 25.21 | 72 | 90 | 41.88 | 4 | 95 | 55.77 | 57.02 | 44.85 | |
| +FastCAOTE | 20.13 | 40.71 | 48.35 | 48.62 | 41.04 | 24.38 | 32.19 | 23.04 | 25.20 | 72 | 90 | 42.33 | 3.5 | 95 | 55.77 | 57.03 | 44.96 | |
| Single Doc. QA | Multi Doc. QA | Summarization | Fewshot Learning | Synthetic | Code | |||||||||||||
| Narrative QA | Qasper | MF-en | HotpotQA | 2WikiMQA | Musique | GovReport | QMSum | MultiNews | TREC | TriviaQA | SAMSum | PCount | PR-en | Lcc | RB-P | Avg. | ||
| Qwen 2.5-7B | 15.75 | 16.94 | 32.38 | 11.89 | 11.88 | 7.95 | 34.33 | 19.91 | 22.67 | 65.5 | 87.05 | 44.75 | 4.22 | 93.08 | 57.74 | 61.84 | 36.74 | |
| 6k | Sink | 7.37 | 16.61 | 25.73 | 11.29 | 11.27 | 5.69 | 31.47 | 18.72 | 22.86 | 64.5 | 84.86 | 44.47 | 3.59 | 41.48 | 55.89 | 55.99 | 31.36 |
| H2O | 3.34 | 14.79 | 23.94 | 11.45 | 11.3 | 5.52 | 14.63 | 14.27 | 22.06 | 55.75 | 51.99 | 28.01 | 1.39 | 9.41 | 54.68 | 38.32 | 22.55 | |
| + CAOTE | 4.78 | 18.06 | 32.49 | 16.23 | 17.28 | 9.57 | 29.81 | 18.04 | 22.86 | 59.5 | 63.05 | 36.91 | 2.7 | 28.25 | 55.13 | 42.42 | 28.57 | |
| + FastCAOTE | 5.69 | 16.99 | 32.62 | 18.22 | 16.58 | 10.48 | 30.3 | 17.71 | 22.88 | 59.5 | 62.95 | 36.29 | 2.1 | 27.65 | 56.3 | 40.65 | 28.56 | |
| TOVA | 15.77 | 15.33 | 30.31 | 19.3 | 13.78 | 9.11 | 30.4 | 19.95 | 22.91 | 61.5 | 83.47 | 42.9 | 1.15 | 21.775 | 57.68 | 57.99 | 31.46 | |
| + CAOTE | 15.81 | 16.07 | 29.39 | 19.4 | 14.15 | 10.8 | 30.89 | 20.54 | 22.86 | 62 | 84.92 | 43.19 | 2.17 | 30 | 57.76 | 57.53 | 32.34 | |
| + FastCAOTE | 15.67 | 16.23 | 30.4 | 19.45 | 13.32 | 10.18 | 30.77 | 20.2 | 22.82 | 61.5 | 83.29 | 43.3 | 1.5 | 28.75 | 57.71 | 58.1 | 32.07 | |
| SnapKV | 14.34 | 16.35 | 31.12 | 17.56 | 14.1 | 8.74 | 31.09 | 20.16 | 22.84 | 60 | 83.8 | 42.99 | 2.91 | 54.17 | 57.48 | 60.26 | 33.62 | |
| + CAOTE | 12.77 | 16.3 | 31.33 | 19.74 | 14.06 | 11.07 | 31.02 | 20.85 | 22.91 | 61.5 | 83.79 | 42.97 | 4.8 | 68.25 | 57.54 | 61.08 | 35.00 | |
| + FastCAOTE | 12.98 | 15.93 | 31.3 | 18.58 | 13.82 | 9.45 | 30.96 | 20.27 | 22.88 | 61.5 | 84.58 | 43.28 | 5.34 | 65.48 | 57.54 | 60.25 | 34.63 | |
| 8k | H2O | 6.1 | 15.55 | 28.29 | 12.37 | 14.65 | 6.24 | 20.78 | 17.22 | 22.44 | 59 | 58.74 | 33.05 | 1.82 | 15.73 | 55.63 | 44.56 | 25.76 |
| + CAOTE | 8.65 | 15.59 | 34.92 | 20.41 | 15.95 | 13.6 | 32.11 | 20.05 | 22.82 | 63.5 | 78.2 | 40.66 | 3.85 | 46.33 | 57.19 | 51.7 | 32.85 | |
| + FastCAOTE | 6.76 | 15.88 | 34.3 | 20.75 | 16.2 | 16.82 | 31.95 | 20.67 | 22.81 | 62.5 | 77.33 | 41.02 | 3.03 | 47.08 | 57.23 | 50.48 | 32.8 | |
| TOVA | 15.69 | 15.55 | 33.09 | 18.37 | 13.99 | 11.26 | 31.33 | 20.17 | 22.82 | 62 | 84.49 | 43.01 | 2.78 | 30.33 | 57.45 | 58.96 | 32.58 | |
| + CAOTE | 16.38 | 15.46 | 32.16 | 17.86 | 14.24 | 12.76 | 31.34 | 20.2 | 22.8 | 61 | 83.97 | 43.23 | 2.01 | 38.83 | 57.45 | 59.41 | 33.07 | |
| + FastCAOTE | 17.06 | 15.55 | 32.32 | 17.57 | 14.14 | 13.03 | 31.43 | 20.3 | 22.78 | 62 | 84.86 | 43.3 | 2.41 | 41.58 | 57.41 | 58.93 | 33.42 | |
| SnapKV | 15.6 | 15.81 | 33.47 | 18.02 | 14.49 | 10.53 | 31.99 | 20.09 | 22.84 | 61 | 84.08 | 43.01 | 4.58 | 64.25 | 57.46 | 60.59 | 34.86 | |
| + CAOTE | 15.55 | 15.57 | 33.89 | 21.08 | 14.43 | 12.38 | 31.41 | 20.73 | 22.83 | 61.5 | 85.11 | 43.39 | 5.22 | 75.75 | 57.44 | 60.35 | 36.04 | |
| + FastCAOTE | 13.38 | 15.77 | 33.97 | 19.78 | 15.08 | 13.01 | 31.44 | 20.69 | 22.77 | 62 | 85.66 | 43.69 | 4.24 | 75.33 | 57.44 | 60.04 | 35.89 | |
| Qwen 2.5-3B | 18.08 | 22.49 | 39.72 | 27.86 | 20.45 | 18.93 | 32.8 | 23.74 | 24.89 | 67.5 | 85.05 | 43.88 | 5 | 40.97 | 51.91 | 47.53 | 35.68 | |
| 6k | Sink | 13.01 | 20.03 | 32.59 | 18.62 | 15.77 | 9.37 | 30.98 | 20.7 | 24.97 | 66.5 | 75.39 | 42.77 | 4 | 14.92 | 52.32 | 50.35 | 30.77 |
| H2O | 5.52 | 18.62 | 27.93 | 12.61 | 15.07 | 4.26 | 14.92 | 13.89 | 24.21 | 58 | 45.94 | 24.93 | 2.91 | 9.1 | 49.5 | 34.54 | 22.62 | |
| + CAOTE | 8.23 | 21.34 | 36.28 | 22.43 | 17.92 | 13.53 | 31.57 | 21.2 | 24.91 | 65 | 74.3 | 39.09 | 4.58 | 21.25 | 50.47 | 42.71 | 30.93 | |
| + FastCAOTE | 9.29 | 20.47 | 35.8 | 21.67 | 18.14 | 13.65 | 31.34 | 20.52 | 24.82 | 64.5 | 75.88 | 39.16 | 5.72 | 20.42 | 50.6 | 44.38 | 31.02 | |
| TOVA | 13.62 | 19.56 | 34.64 | 21.67 | 16.25 | 8.47 | 30.17 | 23.1 | 24.94 | 63.5 | 81.88 | 42.97 | 1.16 | 10.58 | 51.3 | 47.7 | 30.72 | |
| + CAOTE | 13.28 | 19.82 | 35.25 | 22.6 | 15.5 | 8.97 | 30.4 | 23.17 | 24.84 | 64.5 | 81.68 | 43.46 | 2.07 | 13.21 | 50.56 | 47.05 | 31.02 | |
| + FastCAOTE | 13.34 | 19.71 | 35.58 | 22.02 | 15.65 | 8.76 | 30.49 | 23.26 | 24.82 | 65 | 80.92 | 43.66 | 1.75 | 13 | 51.45 | 47.74 | 31.07 | |
| SnapKV | 14.16 | 20.09 | 36.15 | 19.14 | 15.59 | 12.7 | 30.35 | 22.75 | 24.91 | 65 | 83.92 | 43.52 | 5.00 | 32.2 | 51.04 | 47.49 | 32.75 | |
| +CAOTE | 14.3 | 20.13 | 34.89 | 20.32 | 15.06 | 12.85 | 30.61 | 23.16 | 24.9 | 66.5 | 84.75 | 43.3 | 4.75 | 33.9 | 51.48 | 48.31 | 33.08 | |
| +FastCAOTE | 14.33 | 19.48 | 34.59 | 20.8 | 16.54 | 11.85 | 30.68 | 23.19 | 24.93 | 66.5 | 83.54 | 43.55 | 4.62 | 33 | 51.06 | 47.93 | 32.91 | |
| 8k | H2O | 6.16 | 19.84 | 32.32 | 16.01 | 17.74 | 4.99 | 20.21 | 16.49 | 24.54 | 64 | 56.1 | 32.56 | 3.13 | 11.61 | 50.61 | 38.8 | 25.94 |
| +CAOTE | 11.53 | 21.59 | 38.02 | 25.62 | 20.19 | 15.11 | 32.18 | 21.82 | 24.81 | 67.5 | 79.32 | 41.2 | 5.15 | 25.5 | 50.72 | 45.27 | 32.85 | |
| +FastCAOTE | 11.65 | 21.2 | 37.92 | 24.47 | 20.32 | 13.25 | 32.11 | 21.75 | 24.84 | 67.5 | 78.07 | 40.27 | 5.17 | 25.21 | 50.17 | 46.37 | 32.52 | |
| TOVA | 14.66 | 20.93 | 37.77 | 22.57 | 17.08 | 9.63 | 31.12 | 23.17 | 24.83 | 67 | 84.11 | 43.55 | 2.06 | 13.08 | 51.32 | 47.64 | 31.91 | |
| +CAOTE | 13.98 | 21 | 36.91 | 22.97 | 17.05 | 9.93 | 31.25 | 23.45 | 24.9 | 66.5 | 84.14 | 43.86 | 3.07 | 13.92 | 51.14 | 47.883 | 32.00 | |
| +FastCAOTE | 14.84 | 20.66 | 37.45 | 23.05 | 17.07 | 10.16 | 31.2 | 23.47 | 24.85 | 66.5 | 84.01 | 43.41 | 3.31 | 13.25 | 51.19 | 48.11 | 32.03 | |
| SnapKV | 12.76 | 20.88 | 37.1 | 22.49 | 18.19 | 13.83 | 31.33 | 23.37 | 24.8 | 65.5 | 84.88 | 44.49 | 5.2 | 35.83 | 51.31 | 47.82 | 33.74 | |
| +CAOTE | 13.66 | 20.41 | 38.08 | 24.76 | 17.31 | 13.21 | 31.3 | 23.62 | 24.82 | 66.5 | 84.88 | 44.16 | 5.17 | 36.58 | 51.24 | 47.94 | 33.98 | |
| +FastCAOTE | 14.56 | 20.99 | 37.61 | 25.56 | 18.02 | 13.89 | 31.37 | 23.27 | 24.83 | 66.5 | 84.88 | 44.01 | 5 | 35.92 | 51.13 | 48.14 | 34.11 | |
LongBench result for k and K for Llama 3.2-3B-Instruct/3.1-8B-Instruct and Qwen 2.5-3B-Instruct/8B-Instruct are shown in Table Table 7, Table 8 respectively. We also include Sink attention results (Xiao et al., 2024) with budget of k and k. For Llama 3.1-8B-Instruct, TOVA-FastCAOTE performs best for k budget, while SnapKV-FastCAOTE for k budget. For Llama 3.2-3B-Instruct SnapKV-FastCAOTE performs best for both k and k budget. On the other hand, for Qwen 2.5-7B-Instruct, SnapKV-CAOTE performs the best for both k and k, and for Qwen 2.5-3B-Instruct, SnapKV-COATE performs best for k budget, and SnapKV-FastCAOTE performs best for k budget. Additionally, note that all baseline token eviction methods achieve boost in accuracy when using CAOTE or FastCAOTE.
C.2 Perplexity
| Budget | H2O | TOVA | SnapKV | ||||||
| +CAOTE | +FastCAOTE | +CAOTE | +FastCAOTE | +CAOTE | +FastCAOTE | ||||
| Qwen 2.5-7B-Instruct | |||||||||
| 2k | 0.4253 | 0.4422 | 0.3917 | 0.0567 | 0.059 | 0.1077 | 0.0369 | 0.0987 | 0.0307 |
| Qwen 2.5-3B-Instruct | |||||||||
| 2k | 0.2585 | 0.2168 | 0.2154 | 0.0603 | 0.0513 | 0.0507 | 0.0278 | 0.0199 | 0.0196 |
We show perplexity results for Qwen 2.5 models in Table 9 for budget of k. SnapKV-FastCAOTE performs best for both Qwen 2.5-3B-Instruct and 2.5-7B-Instruct, and using CAOTE, all methods achieve improved perplexity.
C.3 Needle in a Haystack
| Budget | H2O | TOVA | SnapKV | ||||||
| +CAOTE | +FastCAOTE | +CAOTE | +FastCAOTE | +CAOTE | +FastCAOTE | ||||
| Qwen 2.5-7B-Instruct | |||||||||
| 6k | 0.206 | 0.312 | 0.3 | 0.292 | 0.292 | 0.286 | 0.32 | 0.33 | 0.332 |
| Qwen 2.5-3B-Instruct | |||||||||
| 6k | 0.212 | 0.288 | 0.27 | 0.282 | 0.286 | 0.288 | 0.304 | 0.324 | 0.336 |
Table 10 shows Needle-in-haystack results for Qwen2.5 models for budget=k. SnapKV-FastCAOTE performs best for both Qwen 2.5-3B-Instruct and 2.5-7B-Instruct, and using CAOTE, all methods achieve improved accurracy.

### v4

@@
 Any norm can be used for computing CAOTE score and based on empirical results we choose -norm.
+Evicting multiple tokens using CAOTE formulation is discussed in Appendix B.
 3.2 CAOTE with general score-based eviction methods
@@
 CAOTE for H2O
-We consider H2O (Zhang et al., 2024b), where the scores () are based on the sum of previous attention scores, leading to during generation-phase as proved in Theorem 24 in Section B.1.
+We consider H2O (Zhang et al., 2024b), where the scores () are based on the sum of previous attention scores, leading to during generation-phase as proved in Theorem 36 in Section D.1.
 In this case, simply dividing each token score by the sum of all scores maps the scores to the range and ensures that new scores follow .
@@
 3.3 Fast CAOTE Computation
-We also propose a compute-efficient version of CAOTE, FastCAOTE, with negligible performance degradation and reduced computation by an order of , where is the hidden dimension of the model.
+We also propose a compute-efficient version of CAOTE, FastCAOTE, with negligible performance degradation and minimal compute overhead relative to CAOTE.
 Here, the pre-eviction attention output () is replaced with mean of values while everything else remains same, that is, CAOTE score for token is:
+We find empirically, that FastCAOTE is highly correlated to CAOTE by observing high () Spearman’s correlation between them for each layer as shown in Table 8.
+We provide theoretical inference time latency for using CAOTE and FastCAOTE in Appendix C, where we observe that FastCAOTE leads to minimal increase in latency.
 4 Results
@@
 We present the accuracy of Llama 3.1-8B-Instruct, Llama 3.2-3B-Instruct and, Qwen 2.5-3B-Instruct, Qwen 2.5-7B-Instruct using baseline eviction methods with budget of k, k, both with and without CAOTE in Table 2 and Table 3.
-We observe that the best performance is given by SnapKV-FastCAOTE for the Llama3 models, while for Qwen 2.5 models SnapKV-CAOTE performs the best.
+We observe that the best average performance is given by SnapKV-FastCAOTE for the Llama3 models, while for Qwen 2.5 models SnapKV-CAOTE performs the best.
 H2O shows improvement with CAOTE, while TOVA, SnapKV also show overall improvements, making their average accuracy closer to dense accuracy.
-Additional results for the k, k budget are shown in Table 7, Table 8 for Llama3 and Qwen 2.5 respectively, in Section C.1, which follow a trend similar to the k, k budgets.
+Additionally, for some QA tasks (Qasper, MF-en, Musique, 2WikiMQA), H2O-(Fast)CAOTE performs best.
+Additional results for the k, k budget are shown in Table 9, Table 10 for Llama3 and Qwen 2.5 respectively, in Section E.1, which follow a trend similar to the k, k budgets.
 4.3 Perplexity
@@
 TOVA-FastCAOTE, SnapKV-CAOTE, and SnapKV-FastCAOTE perform best for k, k, k budgets, respectively, for Llama 3.1-8B-Instruct; for Llama 3.2-3B-Instruct, TOVA-FastCAOTE performs best with k and k budgets and SnapKV-FastCAOTE beats other methods using k and k.
-Perplexity results for Qwen 2.5 models are shown in Table 9 in Section C.2.
+Perplexity results for Qwen 2.5 models are shown in Table 11 in Section E.2.
 4.4 Needle In A HayStack
@@
 We can see in Figure 2 that CAOTE improves the state-of-the-art eviction method and is able to predict beyond their budget constraints.
-Results for Qwen 2.5 models are shown in Table 10 in Section C.3.
-4.5 Ablations
-Difference in Attention Outputs
-In Figure 3, we check the average change in attention score in different layers for various eviction methods to provide insight into the performance gain realized by CAOTE.
-We measure the difference between the dense attention-output and eviction method based attention output, both with and without CAOTE.
-We repeat this for each layer and average the result to verify that CAOTE results in smaller distance to dense attention-output.
-Figure 3 shows the normalized mean squared error between the dense attention output and eviction-based attention output for H2O, TOVA and SnapKV, both with and without CAOTE in dashed-lines and solid-lines respectively.
-We observe that the CAOTE variant always produces a smaller gap from dense attention output (no eviction) compared to baselines, supporting our claim that CAOTE evicts the token which minimizes deviation from true attention output.
-Furthermore, we show that discrepancy in attention output due to eviction (eviction error) causes discrepancy in logits in B.2; to emphasize that optimizing for eviction error results in less discrepancy in logits and, therefore, less deviation on downstream task accuracy.
-Different Prompt Block Sizes
-| Task | H2O | H2O-CAOTE | TOVA | TOVA-CAOTE | SnapKV | SnapKV-CAOTE | ||||||||||||
-| block-size | 32 | 128 | 32 | 128 | 32 | 128 | 32 | 128 | 32 | 128 | 32 | 128 | ||||||
-| HotPotQA∗ | 20.43 | 24.49 | 35.71 | 32.37 | 40.28 | 46.37 | 44.36 | 45.55 | 52.93 | 44.72 | 45.25 | 52.94 | 45.99 | 46.7 | 53.67 | 44.66 | 45.64 | 52.67 |
-| Qasper | 30.8 | 31.94 | 32.58 | 36.74 | 40.79 | 38.93 | 38.65 | 38.83 | 40.45 | 39.39 | 39.79 | 39.90 | 38.71 | 39.22 | 39.80 | 40.18 | 39.74 | 39.69 |
-| GovReport | 5.52 | 5.44 | 7.07 | 30.10 | 30.82 | 31.48 | 29.32 | 29.41 | 31.25 | 29.24 | 29.39 | 31.15 | 29.62 | 29.79 | 31.82 | 30.15 | 30.07 | 32.42 |
-| Trec | 46.50 | 53.00 | 56.00 | 59.00 | 65.5 | 72.00 | 62.00 | 64.00 | 71.50 | 64.00 | 63.00 | 71.50 | 64.00 | 65.50 | 71.50 | 67.50 | 67.50 | 71.50 |
-| LCC | 51.59 | 51.23 | 51.51 | 54.06 | 55.32 | 54.51 | 56.23 | 55.64 | 56.34 | 56.18 | 55.68 | 56.39 | 56.24 | 55.74 | 56.56 | 56.27 | 55.73 | 56.59 |
-Prompt block size dictates how many tokens will be evicted at once.
-A moderately large prompt block size results in faster prompt processing but evicts more tokens each time, while a smaller prompt block size results in slower prefilling.
-To evaluate the impact of prompt block size, we evaluated a subset of LongBench tasks with block prompt sizes of , 128 and , i.e., process the entire prompt as a single large block and display the results in Table 6 for Llama3.2 with budget=k.
-Based on the CAOTE formulation (single token eviction), the smaller the block-size the better CAOTE would perform as seen for block-size= in Table 6, however, even for large prompt block size, CAOTE performs competitive to baseline if not better.
-Note that the infinite prompt block size initially violates the memory/budget constraint, and additionally, we observe out-of-memory errors for several samples for the HotPotQA and GovReport tasks when using infinite block size.
-This problem is not present with block-size=.
+Results for Qwen 2.5 models are shown in Table 12 in Section E.3.
 5 Related Work
@@
 Due to this reason CAOTE’s performance might further improve with smaller prompt filling block-size.
-Appendix B Additional Proofs
-B.1 H2O scores
-Theorem B.1.
+Appendix B Extending CAOTE to Multi-token Eviction
+We start with simple example, without loss of generality consider evicting two tokens in the order . From (8) single token eviction error for token is
+After evicting token , the updated attention output becomes:
+Now, evicting token from updated attention output above yields
+This leads to the following insights:
+The updated attention output can be expressed as
+(28) -
+Eviction error of removing token after removing token is
+(29) -
+Substituting the expressions above, we obtain a closed-form joint eviction score:
+This formulation generalizes to evicting tokens jointly (assuming first tokens here without loss of generality)
+This expression highlights the combinatorial nature of multi-token eviction: for tokens with -token eviction, there are possible combinations.
+This makes exact computation intractable for large , motivating the need for approximate or greedy strategies.
+Importantly, existing token eviction methods typically rely only on attention scores and assume independence between tokens.
+As a result, the relative ranking of tokens remains unchanged even after evictions, which limits their effectiveness in multi-token settings.
+In contrast, CAOTE explicitly models the interplay between attention scores and value vectors, capturing how evicting one token affects the contribution of others.
+This introduces interdependencies between tokens during eviction, which are critical for accurate multi-token decisions.
+Appendix C Inference Latency
+We analyze the computational overhead of CAOTE and FastCAOTE during both prefill and generation phases. The overhead is minimal, especially for FastCAOTE.
+Let be sequence length, the hidden dimension, the KV hidden dimension, the intermediate dimension, the vocabulary size, and the number of hidden layers.
+The total floating-point Operation count are as follows:
+We compute the relative overhead for Llama3.1-8B model with different sequence lengths and show the ratio between prefill (and generation) flops to CAOTE flops in Table 6, 7
+| Sequence Length | Ratio | |
+| with CAOTE | with FastCAOTE | |
+| 4k | 1.6e-4 | 8.9e-5 |
+| 8k | 9.25e-5 | 5.28e-5 |
+| 32k | 6.45e-5 | 3.69e-5 |
+| Sequence Length | Ratio | |
+| with CAOTE | with FastCAOTE | |
+| 512 | 0.08 | 0.046 |
+| 1024 | 0.15 | 0.087 |
+Appendix D Additional Proofs
+D.1 H2O scores
+Theorem D.1.
 Given H2O scores for tokens as , the summation of all , is greater than
@@
 In general for a token , the H2O score is defined as
-B.2 Relation of CAOTE score (eviction error) to logits
+D.2 Relation of CAOTE score (eviction error) to logits
 We show that evicting token based on CAOTE score can lead to smaller discrepancy in final logits which affect the downstream performance.
@@
 Note that for multiple layers, each layer would have its own eviction error which will keep compounding; however, this computing the exact compounded error is non-trivial due to the presence of output layer-norms.
-Appendix C Additional Results
-C.1 LongBench
+Appendix E Additional Results
+| Layer | 1 | 2 | 3 | 4 | 5 | 6 | 7 | 8 | 9 | 10 | 11 | 12 | 13 | 14 |
+| 0.99 | 0.98 | 0.91 | 0.97 | 0.92 | 0.88 | 0.85 | 0.82 | 0.81 | 0.90 | 0.84 | 0.92 | 0.83 | 0.86 |
+| layer | 15 | 16 | 17 | 18 | 19 | 20 | 21 | 22 | 23 | 24 | 25 | 26 | 27 | 28 |
+| 0.91 | 0.88 | 0.94 | 0.96 | 0.92 | 0.95 | 0.99 | 0.99 | 0.96 | 0.98 | 0.98 | 0.88 | 0.99 | 0.96 |
+E.1 LongBench
 | Single Doc. QA | Multi Doc. QA | Summarization | Fewshot Learning | Synthetic | Code | |||||||||||||
@@
 Additionally, note that all baseline token eviction methods achieve boost in accuracy when using CAOTE or FastCAOTE.
-C.2 Perplexity
+E.2 Perplexity
 | Budget | H2O | TOVA | SnapKV | ||||||
@@
 SnapKV-FastCAOTE performs best for both Qwen 2.5-3B-Instruct and 2.5-7B-Instruct, and using CAOTE, all methods achieve improved perplexity.
-C.3 Needle in a Haystack
+E.3 Needle in a Haystack
 | Budget | H2O | TOVA | SnapKV | ||||||
@@
 SnapKV-FastCAOTE performs best for both Qwen 2.5-3B-Instruct and 2.5-7B-Instruct, and using CAOTE, all methods achieve improved accurracy.
+E.4 Ablations
+Difference in Attention Outputs
+In Figure 3, we check the average change in attention score in different layers for various eviction methods to provide insight into the performance gain realized by CAOTE.
+We measure the difference between the dense attention-output and eviction method based attention output, both with and without CAOTE.
+We repeat this for each layer and average the result to verify that CAOTE results in smaller distance to dense attention-output.
+Figure 3 shows the normalized mean squared error between the dense attention output and eviction-based attention output for H2O, TOVA and SnapKV, both with and without CAOTE in dashed-lines and solid-lines respectively.
+We observe that the CAOTE variant always produces a smaller gap from dense attention output (no eviction) compared to baselines, supporting our claim that CAOTE evicts the token which minimizes deviation from true attention output.
+Furthermore, we show that discrepancy in attention output due to eviction (eviction error) causes discrepancy in logits in D.2; to emphasize that optimizing for eviction error results in less discrepancy in logits and, therefore, less deviation on downstream task accuracy.
+Different Prompt Block Sizes
+| Task | H2O | H2O-CAOTE | TOVA | TOVA-CAOTE | SnapKV | SnapKV-CAOTE | ||||||||||||
+| block-size | 32 | 128 | 32 | 128 | 32 | 128 | 32 | 128 | 32 | 128 | 32 | 128 | ||||||
+| HotPotQA∗ | 20.43 | 24.49 | 35.71 | 32.37 | 40.28 | 46.37 | 44.36 | 45.55 | 52.93 | 44.72 | 45.25 | 52.94 | 45.99 | 46.7 | 53.67 | 44.66 | 45.64 | 52.67 |
+| Qasper | 30.8 | 31.94 | 32.58 | 36.74 | 40.79 | 38.93 | 38.65 | 38.83 | 40.45 | 39.39 | 39.79 | 39.90 | 38.71 | 39.22 | 39.80 | 40.18 | 39.74 | 39.69 |
+| GovReport | 5.52 | 5.44 | 7.07 | 30.10 | 30.82 | 31.48 | 29.32 | 29.41 | 31.25 | 29.24 | 29.39 | 31.15 | 29.62 | 29.79 | 31.82 | 30.15 | 30.07 | 32.42 |
+| Trec | 46.50 | 53.00 | 56.00 | 59.00 | 65.5 | 72.00 | 62.00 | 64.00 | 71.50 | 64.00 | 63.00 | 71.50 | 64.00 | 65.50 | 71.50 | 67.50 | 67.50 | 71.50 |
+| LCC | 51.59 | 51.23 | 51.51 | 54.06 | 55.32 | 54.51 | 56.23 | 55.64 | 56.34 | 56.18 | 55.68 | 56.39 | 56.24 | 55.74 | 56.56 | 56.27 | 55.73 | 56.59 |
+Prompt block size dictates how many tokens will be evicted at once.
+A moderately large prompt block size results in faster prompt processing but evicts more tokens each time, while a smaller prompt block size results in slower prefilling.
+To evaluate the impact of prompt block size, we evaluated a subset of LongBench tasks with block prompt sizes of , 128 and , i.e., process the entire prompt as a single large block and display the results in Table 13 for Llama3.2 with budget=k.
+Based on the CAOTE formulation (single token eviction), the smaller the block-size the better CAOTE would perform as seen for block-size= in Table 13, however, even for large prompt block size, CAOTE performs competitive to baseline if not better.
+Note that the infinite prompt block size initially violates the memory/budget constraint, and additionally, we observe out-of-memory errors for several samples for the HotPotQA and GovReport tasks when using infinite block size.
+This problem is not present with block-size=.